OpenStudy (anonymous):

A small rural town has experienced a population decrease at a rate of 25 people every 5 years since 1970. If the population of the town in 1970 was 250, which of the following equations represents the town’s population as a function of time? (Assume 1970 is represented by x = 0, 1971 is represented by x = 1, and so on).

OpenStudy (austinl):

Do you have choices? So I can see what the crap it is supposed to look like?

OpenStudy (anonymous):

A. y = –5x + 250 B. y = –25x + 250 C. –5x + y = 250 D. –x – 25y = 250

OpenStudy (austinl):

Oh dear.

OpenStudy (anonymous):

^ My reaction

OpenStudy (austinl):

Okay, I have a thought... but how to prove it... lmao

OpenStudy (anonymous):

*sighs*

OpenStudy (austinl):

I am leaning towards B....

OpenStudy (anonymous):

Ill take it, I was thinking so as well.

OpenStudy (austinl):

Okies, good work team lol